OpenStudy (anonymous):

Megan is four years younger than her aunt, and the sum of their ages is at least 40. Which inequality would you use to solve determine the ages of Megan and her aunt? A. a - 4 + a ≤ 40 B. a - 4 = 40 C. 2a - 4 ≥ 40 D. m + a - 4 > 40 Somebody help plz

OpenStudy (anonymous):

@princeharryyy

OpenStudy (anonymous):

@PyroYolka

OpenStudy (princeharryyy):

2a - 4 ≥ 40

OpenStudy (anonymous):

let a be the age of aunt then age of megan = a-4 sum of their age = a+a-4 = 2a - 4 as their age's sum is ATLEAST 40 \[2a-4 \ge 40\]

OpenStudy (princeharryyy):

@jada___

OpenStudy (anonymous):

yes ? @princeharryyy

OpenStudy (princeharryyy):

nothing